Use the Garman-Kohlhagen model to calculate the premium of a European US$ put with a strike price of 90 Yen given

Spot exchange rate: 95 Yen per dolar

Volatility: 40 percent

Time to expire: 1 year

US interest rate: 3 percent per annum

Japan interest rate: 6 percent per annum

Show your computations of -d1, -d2, N(-d1) and N(-d2)
Answer #1

The price of a European Put option with strike price K and time of maturity T is given by the formula:

P = K*e^(-rd*t) * N(-d2) - S*e^(-rf*t) *N(-d1)

where d1= ( ln(S/K) + (rd-rf + s^2/2) *t ) / (s*t^0.5)

d2 = ( ln(S/K) + (rd-rf - s^2/2) *t ) / (s*t^0.5) = d1 -s*t^0.5

K is the strike exchange rate = 90 Yen/$

rd is the domestic interest rate =3% = 0.03

t is the time till maturity in years =1

S is the spot exchage rate = 95 Yen/$

rf is the foreign interest rate = 6% =0.06

s is the volatility of the exchange rate = 40% = 0.4

So, d1= (ln(95/90)+(0.03-0.06+0.4^2/2)*1/ (0.4*1^0.5) = 0.260168

So -d1 = -0.260168

d2 == (ln(95/90)+(0.03-0.06-0.4^2/2)*1/ (0.4*1^0.5) = -0.13983

d2 = d1-s*t^0.5 = 0.260168 - 0.4*1^0.5 = -0.13983

So, -d2= 0.13983

N(-d1) = N(-0.260168) = area under normal distribution curve below -0.260168 = 0.397367

N(-d2) = N(0.13983) = area under normal distribution curve below 0.13983 =0.555604

So,

P = K*e^(-rd*t) * N(-d2) - S*e^(-rf*t) *N(-d1)

=90*e^(-0.03*1)* 0.555604 - 95*e^(-0.06*1)*0.397367

=12.9750 yen/$ is the required premium of the European Put option